Can someone please explain how to use this chart or how to get the number from it. There's so many numbers so I don't know what's going on

5.00 3.00 1.00 1.60 2 ?/ 1.40 1.30 1,20 1.15 1501 40 0.85 0.80 0.90 0.70 0.95 0.60 8 0.50 0.40 0.30 0.0 0 0.2 0.3 0.4 0.6 0.7 0.8 0.9 1.0 Reduced pressure, pA
Answer #1

The dimensionless compressibility factor, Z, for a gaseous species is defined as the ratio Z = pv RT. If the gas behaves ideally then Z = 1. The extent to which Z differs from 1 is a measure of the extent to which the gas is behaving non-ideally. The compressibility can be determined from experimental data where Z is plotted versus a dimensionless reduced pressure PR and reduced temperature TR, defined as

PR = P/Pc and TR = T/Tc

Where Pc and Tc denote the critical pressure and temperature, respectively.

The solid lines represent reduced temperature TR values and bottom axis represents the reduced pressure PR values. Draw a vertical line from the calculated or given value of PR where it cuts the solid lines TR, draw horizontal line to get corresponding value of Z.

Looking more closely at the generalized compressibility chart we notice that there are not only TR lines, but also VR' lines (dotted). VR' is called the pseudo-reduced specific volume, since it is defined as

VR values has been seen when PR or TR value are not known followed by similar way of looking to the chart

It can be seen from above chart that the value of Z tends to unity for all temperatures as pressure approach zero and Z also approaches unity for all pressure at very high temperature.